OpenStudy (anonymous):

sqrt of 3x times sqrt15x

OpenStudy (jdoe0001):

\(\bf \sqrt{3}\cdot \sqrt{15}\qquad {\color{brown}{ 15\to 3\cdot 5}}\quad thus \\ \quad \\ \sqrt{3}\cdot \sqrt{15}\implies \sqrt{3}\cdot \sqrt{{\color{brown}{ 3\cdot 5}}}\implies \sqrt{3}\cdot \sqrt{3}\cdot \sqrt{5}\implies ?\)